Use the following corn futures quotes:

  Corn 5,000 bushels
  Contract Month Open High Low Settle Chg Open Int
  Mar 455.125 457.000 451.750 452.000 −2.750 597,913
  May 467.000 468.000 463.000 463.250 −2.750 137,547
  July 477.000 477.500 472.500 473.000 −2.000 153,164
  Sep 475.000 475.500 471.750 472.250 −2.000 29,258

a. How many of the May contracts are currently open?

b. How many of these contracts should you sell if you wish to deliver 110,000 bushels of corn in May?

c. If you actually make delivery, how much will you receive? Assume you locked in the settle price. (Do not round intermediate calculations. Round your answer to the nearest whole number.)
Answer #1

The solution to the given question is provided below:
1.
137547 contracts

2.
=110000/5000=22 contracts

3.
=110000*463.250/100=$509,575.00